Conditional Probability/Bayes' Formula/Independent Events
Not sure which one the below is; can anyone help? thank you!!
A legislative committee consists of 7 Democrats, 5 Republicans, and 1 Independent. A delegation of 3 is selected to visit a small Pacific island republic:
a) How many different delegations are possible?
b) How many different delegations consist of at least 1 Republican?

Answer by stanbon(48558)

A legislative committee consists of 7 Democrats, 5 Republicans, and 1 Independent. A delegation of 3 is selected to visit a small Pacific island republic:
a) How many different delegations are possible?
Ans: 13C3 = (13*12*11)/(1*2*3) = 286
-------------------------------------------------------
b) How many different delegations consist of at least 1 Republican?
Ans: 286 - # of committees with no Repoblican
= 286 - 8C3 = 230
======================================================
Cheers,
Stan H.